OpenStudy (anonymous):

The length of the base of a triangle is 5.3 cm and its height is 6.7 cm. If the area of the triangle is half of the product of its base length and height, the area of the triangle is __________ cm2. Roundto the appropriate significant digits

OpenStudy (anonymous):

17.755.

OpenStudy (solomonzelman):

@CallMeDon, dude! Don't just give the answers, show your work! I won't report, but I am sure others will!

OpenStudy (anonymous):

5.3 x 6.7 = 35.51. Then divide by two.

OpenStudy (anonymous):

Sorry, new to the site. My bad

OpenStudy (shamil98):

A = 1/2 bh \[A= \frac{ 1 }{ 2 } \times 5.3 \times 6.7\] \[A =\frac{35.51 }{ 2 }\]

OpenStudy (solomonzelman):

Cool.... Bye!

OpenStudy (solomonzelman):

@CallMeDon, it's ok just don't do that in the future!

OpenStudy (shamil98):

\[A = 17.7\]

OpenStudy (shamil98):

17.755*

OpenStudy (anonymous):

THANK YOU ALL.YOU GUYS ARE SO LOVE